OpenStudy (anonymous):

I'm studying for a pre-calculus test and I dont really understand Composite functions and how to decompose them. Any help would be great.

hartnn (hartnn):

Composite function means a function of a function. If this makes it more confusing, then look at the following example: h(x) = sin (x^2) this function is made of 2 functions, sine function and square function. so if we want to decompose, we can write let f(x)=sin x, g(x)=x^2. to get f(g(x)), replace all 'x' in f(x) by g(x) and tell me what u get...

OpenStudy (anonymous):

so f(x)=sin x^2?

hartnn (hartnn):

f(x) is sin x, g(x) is x^2 so when u replace x in f(x) by g(x)=x^2 u get f(g(x)) = sin x^2 = h(x) did u get this ? if yes, then try g(f(x))

OpenStudy (anonymous):

wouldn't it be the same thing?

hartnn (hartnn):

nopes, try to put 'x' in g(x) by f(x)=sin x, what do u get ?

OpenStudy (anonymous):

(sin x)^2?

hartnn (hartnn):

yes! thats correct. so are (sinx)^2 and sin (x^2) same ?

OpenStudy (anonymous):

no cause in (sinx)^2 you'd multiply sin and x first then ^2. in sin(x^2) you'd do ^2 first then sin.

hartnn (hartnn):

that is correct, but do u know what a sine function is ?

OpenStudy (anonymous):

not a flippin clue :)

OpenStudy (anonymous):

Right now we're just doing regular functions i guess? If you can even use the word regular in math...

hartnn (hartnn):

no problem, try another one: f(x) = x+1, g(x) = x^3 find, f(g(x)) and g(f(x)) and tell me what u get.

OpenStudy (anonymous):

Alright. f(g(x))=x^3+1 g(f(x))=(x+1)^2

hartnn (hartnn):

did u mistype g(f(x)) ? it should be ^3 , isn't it ?

OpenStudy (anonymous):

OH yeah I just had a brain fart :)

hartnn (hartnn):

then you are correct. now can u find f(g(x)) and g(f(x)) for any functions ?

OpenStudy (anonymous):

I think so, its clearer now. Thank you oodles! :)

hartnn (hartnn):

your welcome :) want another example ?

OpenStudy (anonymous):

No thanks i think i'm good. :) thanks